OpenStudy (anonymous):

fractions: jack spends 1 1/2 as long on his homework as jill, last week, jill spent 9 3/4 hours doing homework how long did jack spend doing homework? the week before,jack spent 23 1/4 hours doing homework, how long did jill spend doing homework? i need the answer right cause then it will restart with the same question but different numbers

OpenStudy (anonymous):

i already gave you answer for a similar question

OpenStudy (anonymous):

shanna, your answer was wrong, i thought it was right (the second answer)

OpenStudy (anonymous):

the first answer as right but the second was wrong

OpenStudy (anonymous):

i need a answer urgently

OpenStudy (anonymous):

for jack y=1 1/2x 1 1/2(9 3/4) what would make it easier is to turn them into improper fractions

OpenStudy (anonymous):

for jill 23 1/4=1 1/2x

OpenStudy (anonymous):

all i need is a answer,

OpenStudy (anonymous):

no need for explaining, just to say

OpenStudy (anonymous):

cause the i dont have time for the explaining right now, as i struggled for 2 day cause of the second question

OpenStudy (anonymous):

jack= 117/8 jill= 31/2 ...i think

OpenStudy (anonymous):

i hate my maths teacher sometimes, or hate my self for not understanding it, >:(

OpenStudy (anonymous):

thank you i will hope it is right

OpenStudy (anonymous):

but im gonna wait for the other person to reply his answer

OpenStudy (anonymous):

1. 9 3/4 = 39/4 39/4 * 3/2 = 117/8 or 14 5/8 2.93/4 = j * 3/2 = 93/4 / 3/2 = 15 1/2

OpenStudy (anonymous):

amk 11 got the same as me, but im still not sure if the same :/

OpenStudy (anonymous):

im gonna check wait and ill tell if its right

OpenStudy (anonymous):

THANK YOU SO MUCH , NOW I DONT NEED TO WORRY, :D